OpenStudy (anonymous):

If f(1) = 0, what are all the roots of the function f(x)=x^3+3x^2-x-3? Use the Remainder Theorem. x = –1, x = 1, or x = 3 x = –3, x = –1, or x = 1 x = –3 or x = 1 x = –1 or x = 3

OpenStudy (loser66):

your idea?

OpenStudy (anonymous):

Sorry I had to step away from the computer for a while, and I thought: x = –1, x = 1, or x = 3 but it wasn't right.

OpenStudy (loser66):

how to attack it? don't guess, give me your strategy

OpenStudy (anonymous):

I don't understand the Remainder Theorem at all, so I looked to see what all of the coefficients had in common. There's a negative one, a positive one, a negative three, and a positive three, so with x = –1, x = 1, or x = 3 the -1, +1, and +3 are taken care of and you can combine the -1 and the +3 to get your -3. I know this probably makes no sense, but it's what my brain came up with in desperation.

OpenStudy (loser66):

if f(1) =0 that means 1 is one of the roots. So that (x-1) is one of the factor of the expression. use long division to get other

OpenStudy (loser66):

do you know long division or synthetic ?

OpenStudy (anonymous):

I have been taught both but only understand long division.

OpenStudy (loser66):

ok, apply it, show me your work.

OpenStudy (anonymous):

k one minute

OpenStudy (anonymous):

x^3+3x^2-x-3 / x-1 x^2(x-1)= x^3 -x^2 x^3+3x^2 -(x^3 -x^2) =4x^2 4x(x-1)= 4x^2 -4x 4x^2 -x 4x^2 -4x =3x 3(x-1)=3x -3 3x-3 -(3x -3) =0 Answer: x^2 +4x +3

OpenStudy (loser66):

very very good job. so, solve for the quadratic x^2 +4x +3 , what do you have ?

OpenStudy (anonymous):

1, 4, and 3?

OpenStudy (loser66):

4,3 are not the roots of that quadratic, let solve: x^2 +4x +3 = (x+1)(x+3) ok?

OpenStudy (anonymous):

Oh yeah! Sorry, I remember now. That makes sense. Thank you.